Let alphabet Σ = {a, b, c}, and consider L1 = {w ∈ Σ ∗ | more than half the symbols in w are c’s}. Prove that L1 is not FS using the pumping lemma.
Answer #1

The language is L1 = {we {a,b,c}* more than half the symbols in w are c} . To prove that the language is not FS, use the pumping lemma as follows.

Let p be the pumping length. Then consider the word w = a^pc^{p+1} . In this word, more than half the symbols are 'c', hence w = apcP+1E L1 .

Now consider any breakup of the word w = xyz such that:

  • \lvert xy\rvert \leq p
  • \lvert y\rvert \geq 1.

Then as the first p letter in w are 'a's, the entire string y can only consist of 'a's. Hence y = a^k, k \geq 1 .

Now if the language were regular, then for each i \geq 0 , it must be that xy^iz \in L1 .

Consider i = 2 . Then w' = xy^2z = a^{p+k}c^{p+1} , as an additional y will contribute k additional 'a's to the word.

But in w' , there are p+1 c's, while the length of the word is 2p+k+1 \geq 2p+2 as k \geq 1 .

This means that at most half of the symbols are 'c's, thus violating the property for the language L1, hence w' \notin L1 .

This contradics the pumping lemma, hence proving that L1 is not FS.

Comment in case of any doubts.